Please assist with the question highlighted in yellow.

Fitting a straight line to a set of data yields the following prediction line. Complete (a) through (c) below. Y11-0.2X D. The Y-intercept, bo-0.2, implies that when the value of X is 0, the mean value of Y is -0.2. b. Interpret the meaning of the slope, b1. Choose the correct answer below. A. The slope, b0.2, implies that the average value of Y is -0.2 O B. The slope, b1 = 11, implies that for each increase of 1 unit in X, the value of Y is expected to increase by 1 1 units. ° C. The slope, b1-02, implies that for each increase of 1 unit in X, the value of Y is expected to increase by 0.2 units D. The slope, b1-0.2, implies that for each increase of 1 unit in X, the value of Y is estimated to decrease by 0.2 units. C. Predict the mean value of Y for X= 6. Y(Type an integer or a decimal.)
Answer #1

Given regression equation is,

\hat Y_i = 11-0.2X_i

c) We want to predict the mean value of Y for X = 6 by substituting X=6 in above equation we get,

\hat Y_i = 11-(0.2*6)

= 11-1.2

\hat Y_i = 9.8